Session: Nov 2003 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Sergio Poppi at Mon Nov 3 01:55:54 2003) Success :
Everithing ok.

Urumqi (zhanghua at Mon Nov 3 02:45:17 2003) Minor failures :
The cooling receiver is fail,so we changed to use old non-cooling receiver

Shanghai (Xinyong Huang at Mon Nov 3 04:50:29 2003) Success :
We start the schedule at scan no0057, because of the scans No0001 to No0056 are out of our antenna's limit.

Onsala (Michael Lindqvist at Mon Nov 3 07:04:40 2003) Minor failures :
TO HIGH WIND: Antenna stowed at 00:06.

Effelsberg (A. Kraus at Mon Nov 3 08:50:57 2003) Minor failures : WIND HA-E
We stopped because of high winds between UT1814-1940. During some scans, the antenna was partly shadowed by the surrounding hills.

Torun (K.Borkowski at Mon Nov 3 09:36:25 2003) Failure : RECO
Our P&G recorder is still unusable; see the feedback for N03C1.

Jodrell_Bank (Alastair Gunn at Mon Nov 3 10:42:32 2003) Minor failures : POIN
Telescope drive tripped at 14:38; schedule restarted at 14:44. Some lost time on J1924+1540. No other known problems.

Noto (P. Cassaro at Mon Nov 3 13:28:48 2003) Success :

Westerbork (Tony Foley at Tue Nov 4 15:00:35 2003) Success : HA-E
VERY many scans are outside HA range. We stopped at 22:53 as all scans after that were unobservable.

Hartebeesthoek (Jonathan Quick at Wed Nov 5 06:49:34 2003) Success :

Onsala (Michael Lindqvist at Wed Nov 5 07:34:15 2003) Minor failures : RECE
Addition to previous report: Problem with IF2. Bad data from 23:00 (even VC:s).